Check my work Zachary Chicken Corporation processes and packages chicken for grocery stores. It purchases chickens from farmepoints Complete this question by entering your answers in the tabs below. eBook Print Req A1 Reg A1 Reqaz Req A2 Req A3 Req BeBook Complete this question by entering your answers in the tabs below. Print References Req A1 Req A2 Req A3 Req B1 Reg B2eBook Print Complete this question by entering your answers in the tabs below. References Req A1 Req A2 Req A3 Req B1 Req B2eBook Complete this question by entering your answers in the tabs below. Print References Req A1 Req A2 Req A3 Req B1 Req B2points Complete this question by entering your answers in the tabs below. eBook Print References Req A1 Req A2 Req A3 Req B1eBook Print Complete this question by entering your answers in the tabs below. References Req A1 Reg A2 Req A3 Req B1 Req B1eBook Print Complete this question by entering your answers in the tabs below. References Req A1 Req A2 Req A3 Req B1 Req B2
Answer #1

Ans.a-1

Drumsticks Breasts
Total Quantity to be used as weights 2500 Pounds 6000 Pounds
Allocation of Joint Cost =(2282+23000)*2500/8500 =(2282+23000)*6000/8500
Joint Cost 7435.88 17846.12

Ans.a-2

Chicken Drumsticks Chicken Breasts
Revenue 4500 15600
Cost of Goods sold 7435.88 17846.12
Gross Margin -2935.88 -2246.12

Ans.a-3

No, the product should not be discontinued since it is a by-product and cost of further processing is less than the market price after further processing. Hence, it is giving a net benefit of $3828.82 to produce drumsticks.

Ans.b-1

Drumsticks Breasts
Total Market Value to be used as weights 4500 15600
Allocation of Joint Cost =(2282+23000)*4500/20100 =(2282+23000)*15600/20100
Joint Cost 5660.15 19621.85

Ans.b-2

Chicken Drumsticks Chicken Breasts
Revenue 4500 15600
Cost of Goods sold 5660.15 19621.85
Gross Margin -1160.15 -4021.85

Ans.c-1

Additional Sale by producing Chicken Steak 27560
Less: Additional Cost to produce Chicken Steak 3300
Less: Chicken Breasts Sales Foregone 15600
Net Benefit in producing Chicken Steak 8660

Since chicken steak is generating additional profit of $8660, chicken steak should be produced.

Ans.c-2

Additional profit by producing chicken steak is $8,660 (Profit will be increased by $8,660).